# A theoretical analysis on the resolution of parametric PDEs via Neural Networks designed with Strassen algorithm

### v2

@@ 4 A Neural Network for Matrix Invertion
[…]
@@
[…]
@@
Figure 1 visualizes the process that follows an input matrix $A$ in the MNN $\Phi^{\operatorname{AUX}^{i}_{n,\varrho}}_{\varepsilon}$. Basically, the operations that appear in the informal sketch (4.13) are replaced by the approximations performed by the MNNs introduced previously.
Step 4: Proof of (4.7) for $N\geq 2$.
Equations (4.4) and (4.16) give us
@@
$\left\lVert P^{i-1}(A)-\left[R\left(\Phi^{\operatorname{AUX}^{i}_{n,\varrho}}_{\varepsilon}\right)(A)\right]_{n+1:2n,1:n}\right\rVert_{2}\leq\varepsilon$ (4.19)
@@
$=\left\lVert\left(\left(\frac{A}{2}\right)^{2}-R\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\right)\left(\frac{A}{2}\bigg{|}\frac{A}{2}\right)\right)\left(\frac{A+\operatorname{I}_{n}}{2}\right)\right\rVert_{2}$
@@
as well as (4.4) and (2.1).
Let us suppose that the result (4.19) is true for $i\geq 2$ and prove it for $i+1$. Using (4.18) and the inductive hypothesis (4.19), we get that
@@
which implies (4.19). This also implies (4.7) because for all $A\in\mathbb{R}^{n\times n}$:
$R\left(\Phi^{\operatorname{NEU}^{N}_{n,\varrho}}_{\varepsilon}\right)(A)$
$=2^{2^{N}-1}R\left(\Phi_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}\right)\left(\operatorname{SQR}^{N-1}_{2^{1-2^{N}}\varepsilon}(A)+2^{-2^{N-1}}\operatorname{I}_{n}\middle|\operatorname{PROD}_{2^{1-2^{N}}\varepsilon}^{N-1}(A)\right)$
$=\operatorname{PROD}_{\varepsilon}^{N}(A).$
@@
Step 5: Proof of (4.8) for $N\geq 2$.
We can rapidly deduce with (4.14) that
@@
Step 6: Proof of (4.9) for $N\geq 2$.
For the number of layers (4.9), we can similarly to Step 5 get that
$L\left(\Phi_{\varepsilon}^{\operatorname{NEU}^{N}_{n,\varrho}}\right)$ $=N\left(L\left(\Phi_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}^{\operatorname{\operatorname{STR}_{n,\varrho}}}\right)+1\right)$
$\leq N\left[2\log_{2}n+5+L\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)\right],$
using again (4.14) and (3.18). ∎
We can finally get a MNN approximating the inverse of some matrices.
Theorem 4.6.
$N(\varepsilon,\delta)=\left\lceil\log_{2}\left(\max\left\{\frac{\log_{2}(\varepsilon(1-\delta))}{\log_{2}\delta},2\right\}\right)\right\rceil.$ (4.20)
for all $A\in\mathbb{R}^{n\times n}$ such that
and satisfying:
$M\left(\Phi^{\operatorname{INV}_{n,\varrho}^{\alpha}}_{\varepsilon,\delta}\right)$ $\leq 14n^{\log_{2}7}\left(N\left(\frac{\varepsilon}{2\alpha},\delta\right)-1\right)\left(M\left(\Pi^{\varrho}_{2^{-2^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}}\frac{\min\left\{\frac{\varepsilon}{\alpha},\frac{1}{4}\right\}}{16n^{3}},2n}\right)+12\right)$
$+n^{2}\left(L\left(\Pi_{2^{-2^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}}\frac{\min\left\{\frac{\varepsilon}{\alpha},\frac{1}{4}\right\}}{16n^{3}},2n}^{\varrho}\right)-14N\left(\frac{\varepsilon}{2\alpha},\delta\right)+20+2\log_{2}n\right)$ (4.22)
$L\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{\alpha}_{n,\varrho}}\right)$ $\leq N\left(\frac{\varepsilon}{2\alpha},\delta\right)\left[2\log_{2}n+5+L\left(\Pi_{2^{-2^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}}\frac{\min\left\{\frac{\varepsilon}{\alpha},\frac{1}{4}\right\}}{16n^{3}},2n}^{\varrho}\right)\right],$ (4.23)
if $N\left(\frac{\varepsilon}{2\alpha},\delta\right)=1$.
Note that $\varepsilon\mapsto N(\varepsilon,\delta)$ decreases with $\varepsilon$ when $\delta\in(0,1)$.
@@
We first focus on proving (4.21).
We know that the identity
@@
$\begin{split}\left\lVert A^{-1}-\alpha\sum_{k=0}^{2^{N}-1}(\operatorname{I}_{n}-\alpha A)^{k}\right\rVert_{2}&=\alpha\left\lVert(\operatorname{I}_{n}-(\operatorname{I}_{n}-\alpha A))^{-1}-\sum_{k=0}^{2^{N}-1}(\operatorname{I}_{n}-\alpha A)^{k}\right\rVert\\ &\leq\alpha\frac{\left\lVert\operatorname{I}_{n}-\alpha A\right\rVert_{2}^{2^{N}}}{1-\left\lVert\operatorname{I}_{n}-\alpha A\right\rVert_{2}}\leq\alpha\frac{\delta^{2^{N}}}{1-\delta}\end{split}$ (4.26)
@@
Our plan is then to use the MNN in Theorem 4.5 to approximate matrix inversion. It is easy to check that for $N(\varepsilon,\delta)$ given by (4.20) we get the inequality:
$\frac{\delta^{2^{N(\varepsilon,\delta)}}}{1-\delta}\leq\varepsilon.$ (4.27)
@@
$M\left(\Phi^{\operatorname{IN}}_{n}\right)=n+n^{2}.$ (4.28)
@@
$\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{\alpha}_{n,\varrho}}\coloneq((\mathcal{L}^{1},B^{1},\alpha^{1}),\dots,(\alpha\mathcal{L}^{L},\alpha B^{L},\alpha^{L}))\circ\Phi_{n,\alpha}^{\operatorname{IN}}.$
@@
In the second estimate, we have used (4.26) and (4.7), and in the third one (4.27).
@@
$M\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{\alpha}_{n,\varrho}}\right)=M\left(\Phi_{\min\left\{\frac{\varepsilon}{2\alpha},\frac{1}{8}\right\}}^{\operatorname{NEU}^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}_{n,\varrho}}\right)+M\left(\Phi^{\operatorname{IN}}_{n,\alpha}\right).$ (4.29)
In fact, if $N\left(\frac{\varepsilon}{2\alpha},\delta\right)\geq 2$ we obtain (4.22) from (4.29), (4.8) and (4.28). Similarly, if $N\left(\frac{\varepsilon}{2\alpha},\delta\right)=1$, we obtain (4.24) from (4.29), (4.10) and (4.28).

### v3

@@ 4 Conclusions
$A\frac{1}{2}A|A$$\frac{1}{2}A|0$$\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},\frac{1}{2}}$$\Phi^{\operatorname{FILL}}_{n,L}$$\Phi^{\operatorname{DUP}}_{n}$$\operatorname{SQR}^{1}_{\varepsilon}(A)$$\operatorname{PROD}^{1}_{\varepsilon}(A)$$\operatorname{SQR}^{1}_{\varepsilon}(A)|\operatorname{SQR}^{1}_{\varepsilon}(A)$$\operatorname{SQR}^{1}_{\varepsilon}(A)+\frac{\operatorname{I}_{n}}{2^{2^{1}}}|\operatorname{PROD}^{1}_{\varepsilon}(A)$$\Phi^{\operatorname{MIX}}_{n,1}$$\Phi_{\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}$$\Phi_{\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}$$\operatorname{SQR}^{2}_{\varepsilon}(A)$$\operatorname{PROD}^{2}_{\varepsilon}(A)$$\operatorname{SQR}^{i-1}_{\varepsilon}(A)|\operatorname{SQR}^{i-1}_{\varepsilon}(A)$
@@
$\Phi_{n,i-1}^{\operatorname{MIX}}$$\Phi_{\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}$$\Phi_{\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}$$\operatorname{SQR}^{i}_{\varepsilon}(A)$$\operatorname{PROD}^{i}_{\varepsilon}(A)\cdots$
@@
Figure 1 visualizes the process that follows an input matrix $A$ in the MNN $\Phi^{\operatorname{AUX}^{i}_{n,\varrho}}_{\varepsilon}$. Basically, the operations that appear in the informal sketch (B.11) are replaced by the approximations performed by the MNNs introduced previously in a systematic way.
Step 4: Proof of (B.5) for $N\geq 2$. Equations (B.2) and (B.15) give us
@@
$\left\lVert P^{i-1}(A)-\left[R\left(\Phi^{\operatorname{AUX}^{i}_{n,\varrho}}_{\varepsilon}\right)(A)\right]_{n+1:2n,1:n}\right\rVert_{2}\leq\varepsilon$ (B.18)
@@
as well as (B.14), (B.2) and (2.1).
Let us suppose that the result (B.18) is true for $i\geq 2$ and prove it for $i+1$. Using (B.17) and the inductive hypothesis (B.18), we get that
@@
which implies (B.18). This also implies (B.5) because for all $A\in\mathbb{R}^{n\times n}$:
$R\left(\Phi^{\operatorname{NEU}^{N}_{n,\varrho}}_{\varepsilon}\right)(A)=2^{2^{N}-1}R\left(\Phi_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}\right)\left(\operatorname{SQR}^{N-1}_{2^{1-2^{N}}\varepsilon}(A)+2^{-2^{N-1}}\operatorname{I}_{n}\middle|\operatorname{PROD}_{2^{1-2^{N}}\varepsilon}^{N-1}(A)\right)=\operatorname{PROD}_{\varepsilon}^{N}(A).$
@@
Step 5: Proof of (B.6) for $N\geq 2$. We can rapidly deduce with (B.12) that
@@
Step 6: Proof of (B.7) for $N\geq 2$. For the number of layers (B.7), we can similarly to Step 5 get that
$L\left(\Phi_{\varepsilon}^{\operatorname{NEU}^{N}_{n,\varrho}}\right)=N\left(L\left(\Phi_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}^{\operatorname{\operatorname{STR}_{n,\varrho}}}\right)+1\right)\leq N\left[2\log_{2}n+5+L\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)\right],$
using again (B.12) and (3.17). ∎
Let us now prove Theorem 3.7:
@@
We first focus on proving (3.21). We know that the identity
@@
$\left\lVert A^{-1}-\alpha\sum_{k=0}^{2^{N}-1}(\operatorname{I}_{n}-\alpha A)^{k}\right\rVert_{2}=\alpha\left\lVert(\operatorname{I}_{n}-(\operatorname{I}_{n}-\alpha A))^{-1}-\sum_{k=0}^{2^{N}-1}(\operatorname{I}_{n}-\alpha A)^{k}\right\rVert\leq\alpha\frac{\delta^{2^{N}}}{1-\delta}$ (B.19)
@@
Our plan is then to use the MNN in Lemma B.3 to approximate matrix inversion. It is easy to check that for $N(\varepsilon,\delta)$ given by (3.20) we get the inequality:
$\frac{\delta^{2^{N(\varepsilon,\delta)}}}{1-\delta}\leq\varepsilon.$ (B.20)
@@
$M\left(\Phi^{\operatorname{IN}}_{n}\right)=n+n^{2}.$ (B.21)
@@
$\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{\alpha}_{n,\varrho}}\coloneq((\mathcal{L}^{1},B^{1},\alpha^{1}),\dots,(\alpha\mathcal{L}^{L},\alpha B^{L},\alpha^{L}))\odot\Phi_{n,\alpha}^{\operatorname{IN}}.$
@@
In the second estimate, we have used (B.19) and (B.5), and in the third one (B.20).
@@
$M\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{\alpha}_{n,\varrho}}\right)=M\left(\Phi_{\min\left\{\frac{\varepsilon}{2\alpha},\frac{1}{8}\right\}}^{\operatorname{NEU}^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}_{n,\varrho}}\right)+M\left(\Phi^{\operatorname{IN}}_{n,\alpha}\right).$ (B.22)
In fact, if $N\left(\frac{\varepsilon}{2\alpha},\delta\right)\geq 2$ we obtain (3.22) from (B.22), (B.6) and (B.21). Similarly, if $N\left(\frac{\varepsilon}{2\alpha},\delta\right)=1$, we obtain (3.24) from (B.22), (B.8) and (B.21).